OpenStudy (anonymous):

1. In the equation E = 5mz, what happens to z if m stays the same and E is multiplied by 7?

OpenStudy (ipwnbunnies):

E is on the left side, so the left side becomes 7*E. We need to balance the equation on the right side now. Since we're keeping 5 and 'm' constant, we'll need to multiply z by 7, also. 7*E = 5*m*(7*z) 7E = 35mz